OpenStudy (anonymous):

implicit differentiation ln(7+e^xy)=y

OpenStudy (anonymous):

This is what im learning right now!

OpenStudy (anonymous):

do you know how to do it

OpenStudy (anonymous):

take derivative of both sides. \[\dfrac{d}{dx}\ln(\space f(x) \space) = \dfrac{f'(x)}{f(x)} \]

OpenStudy (anonymous):

So \(\dfrac{d}{dx}\ln (7+e^{xy}) = \dfrac{\dfrac{d}{dx}(7+e^{xy})}{7+e^{xy}} = y'\)

OpenStudy (anonymous):

Can you handle it now?

OpenStudy (anonymous):

what do I do from there

OpenStudy (queelius):

You'll have to keep applying the chain rule. So, focus on \[\frac{ d }{ dx } (7+e^{xy})\] \[e^{xy} \frac{ d }{ dx } (xy)\] What's the next step? Hint: product rule.